arricake:

Three packages of blueberries costs $7. How many packages of blueberries can you buy for $42?

KPalm:

you have $42and each package is $7 - how many times can 7 fit into 42?

arricake:

@kpalm wrote:
you have $42and each package is $7 - how many times can 7 fit into 42?
idk

KPalm:

In other words, what is 42 divided by 7?

Krissy525:

@arricake wrote:
@kpalm wrote:
you have $42and each package is $7 - how many times can 7 fit into 42?
idk
divde 72/7

Krissy525:

42 not 72

arricake:

18

KPalm:

@arricake wrote:
18
try again!

arricake:

@kpalm wrote:
@arricake wrote:
18
try again!
i give up

Krissy525:

put 42/7 in a calculator

KPalm:

If I'm not mistaken, this site actually has a built in calculator, but not sure if you need to practice long division (are they still doing that these days?)

arricake:

@krissy525 wrote:
put 42/7 in a calculator
i did

Krissy525:

@kpalm wrote:
If I'm not mistaken, this site actually has a built in calculator, but not sure if you need to practice long division (are they still doing that these days?)
yes and there is a calculator

arricake:

@krissy525 wrote:
@kpalm wrote:
If I'm not mistaken, this site actually has a built in calculator, but not sure if you need to practice long division (are they still doing that these days?)
yes and there is a calculator
its 10

Krissy525:

@arricake wrote:
@krissy525 wrote:
@kpalm wrote:
If I'm not mistaken, this site actually has a built in calculator, but not sure if you need to practice long division (are they still doing that these days?)
yes and there is a calculator
its 10
no

KPalm:

@arricake wrote:
@krissy525 wrote:
@kpalm wrote:
If I'm not mistaken, this site actually has a built in calculator, but not sure if you need to practice long division (are they still doing that these days?)
yes and there is a calculator
its 10
try again, it must have been entered incorrectly

Krissy525:

1 attachment
arricake:

@krissy525 wrote:
@arricake wrote:
@krissy525 wrote:
@kpalm wrote:
If I'm not mistaken, this site actually has a built in calculator, but not sure if you need to practice long division (are they still doing that these days?)
yes and there is a calculator
its 10
no
I GIVE UP ALREADY

carmelle:

You can also solve this question by setting up a proportion. \[\frac{ 3~packages }{ $7 }=\frac{ x }{ $42 }\] First, you cross-multiply. \[(7)(x)=(3)(42)\] \[7x = 126\] Divide both sides by 7: \[\frac{ 7x }{ 7 }=\frac{ 126 }{ 7 }\] \[x = 18\] That means you can buy 18 packages of blueberries for $42.

Krissy525:

true, i didnt think of that

@carmelle wrote:
You can also solve this question by setting up a proportion. \[\frac{ 3~packages }{ $7 }=\frac{ x }{ $42 }\] First, you cross-multiply. \[(7)(x)=(3)(42)\] \[7x = 126\] Divide both sides by 7: \[\frac{ 7x }{ 7 }=\frac{ 126 }{ 7 }\] \[x = 18\] That means you can buy 18 packages of blueberries for $42.

Krissy525:

so we had it wrong...

KPalm:

Oh shoot, I was reading it wrong! THREE cost 7! so after dividing take that answer and multiply by 3. The answer to 42/7 is 6, though

Krissy525:

yea same

@kpalm wrote:
Oh shoot, I was reading it wrong! THREE cost 7! so after dividing take that answer and multiply by 3. The answer to 42/7 is 6, though

carmelle:

Next time, please make sure you read the question thoroughly so that you don't confuse the person that is asking the question. :)

KPalm:

@carmelle wrote:
Next time, please make sure you read the question thoroughly so that you don't confuse the person that is asking the question. :)
Yes, thank you! Absolutely my bad there :)

Krissy525:

@carmelle wrote:
Next time, please make sure you read the question thoroughly so that you don't confuse the person that is asking the question. :)
sorry